Job P = $ 68,145 Job Q = predetermined overhead rate * Actual Total machine hours of Job Q Job Q = $ 11.55 Per machine hour *[The following information applies to the questions displayed below.] Sweeten Company had no jobs in progress at the beginnin5. What was the total manufacturing cost assigned to Job Q? 6. If Job Q included 30 units, what was its unit product cost? 7.
Answer #1

Plantwide overhead rate = (2500*3.2+1500*4+32200)/4000 = 11.55 per hour

5) Total manufacturing cost of Job Q

Job Q
Direct material 17000
Direct labor 14700
Overhead (5300*11.55) 61215
Total manufacturing cost 92915

6) Unit product cost = 92915/30 = 3097.17 or 3097

7) Selling price of Job

Job P Job Q
Direct material 31000 17000
Direct labor 35400 14700
Overhead 68145 61215
Total manufacturing cost 134545 92915
Markup 107636 74332
Selling price 242181 167247

8) Cost of goods sold = 134545+92915 = 227460